\begin{thebibliography}{AHRT}
\itemsep=2pt

\bibitem{AR} J. Ad\'{a}mek and J. Rosick\'{y}, {\em Locally Presentable and Accessible Categories}, Cambridge 
University Press 1994. 

\bibitem{B} J. T. Baldwin, {\em Categoricity}, Amer. Math. Soc. 2009. 

\bibitem{Ba} B. Banaschewski, {\em More on compact Hausdorff spaces and finitary duality}, Canad. J. Math. 36
(1984), 1113-1118.

\bibitem{Bn} P. Bankston, {\em Some obstacles to duality in topological algebras}, Canad. J. Math. 34 (1982), 80-90.

\bibitem{J} T. Jech, {\em Set Theory}, Academic Press 1978.

\bibitem{Ki} J. Kirby, {\em Abstract elementary categories}, preprint 2008.

\bibitem{K} D. W. Kueker, {\em Abstract elementary classes and infinitary logic}, Ann. Pure Appl. Logic 156
(2008), 274-286.

\bibitem{L} M. J. Lieberman, {\em Topological and Category-Theoretic Aspects of Abstract Elementary Classes},
Thesis, The University of Michigan 2009.
 
\bibitem{L1} M. J. Lieberman, {\em Category theoretic aspects of abstract elementary classes}, preprint, April 2010.
 
\bibitem{MP} M. Makkai and R. Par\' e, {\em Accessible categories: The Foundation of Categorical Model Theory},
Cont. Math. 104, AMS 1989. 

\bibitem{R} J. Rosick\' y, {\em Categories of models of languages $L_{\kappa,\lambda}$}, In: Abstr. 9th Winter
School on Abstract Analysis, Mathem. Inst. Prague 1981, 153-157.
 
\bibitem{R1} J. Rosick\' y, {\em More on directed colimits of models}, Appl. Cat. Struct. 2 (1994), 71-76.
 
\bibitem{R2} J. Rosick\' y, {\em Accessible categories, saturation and categoricity}, J. Symb. Logic 62 (1997),
891-901.

\bibitem{R3} J. Rosick\' y, {\em Elementary categories}, Arch. Math. 52 (1989), 284-288.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{latexsym}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage[dvips]{graphicx}
\usepackage{xypic}
\input xy
\xyoption{all}

\newtheorem{theo}{Theorem}[section]
\newtheorem{lemma}[theo]{Lemma}
\newtheorem{propo}[theo]{Proposition}
\newtheorem{defi}[theo]{Definition}

\newtheorem{rem}[theo]{Remark}
\newtheorem{exam}[theo]{Example}

\newcommand\Ind{\operatorname{Ind}}

\newcommand\Str{\operatorname{\bf Str}}

\newcommand\rank{\operatorname{rank}}

\newcommand\op{\operatorname{op}}

\newcommand\id{\operatorname{id}}
\newcommand\Id{\operatorname{Id}}

\newcommand\Mod{\operatorname{\bf Mod}}
\newcommand\Elem{\operatorname{\bf Elem}}
\newcommand\Emb{\operatorname{\bf Emb}}
\newcommand\Set{\operatorname{\bf Set}}

\newcommand\Pos{\operatorname{\bf Pos}}

\newcommand\ca{\mathcal {A}}

\newcommand\cc{\mathcal {C}}
\newcommand\cd{\mathcal {D}}

\newcommand\ch{\mathcal {H}}

\newcommand\ck{\mathcal {K}}
\newcommand\cl{\mathcal {L}}
\newcommand\cm{\mathcal {M}}

\date{May 11, 2010}
 
\begin{document}
\title{Abstract elementary classes and accessible categories}
\author[T. Beke and J. Rosick\'{y}]
{T. Beke and J. Rosick\'{y}$^*$}
\thanks{ $^*$ Supported by MSM 0021622409.} 
\address{\newline T. Beke\newline
Department of Mathematics\newline
University of Massachusetts Lowell\newline
One University Avenue, Lowell, MA 01854, U.S.A.\newline
tibor\_beke@uml.edu
\newline J. Rosick\'{y}\newline
Department of Mathematics and Statistics\newline
Masaryk University, Faculty of Sciences\newline
Kotl\'{a}\v{r}sk\'{a} 2, 611 37 Brno, Czech Republic\newline
rosicky@math.muni.cz
}
\begin{abstract}
We compare abstract elementary classes of Shelah with accessible categories having directed colimits. 
\end{abstract}
\keywords{abstract elementary class, accessible category}

\maketitle

\section{Introduction}
Infinitary model theory deals with the properties of models of infinitary first-order theories.  M.~Makkai and R.~Par\' e 
\cite{MP} initiated a category-theoretic investigation of the subject by introducing \emph{accessible categories} 
and \emph{accessible functors}.  Accessible categories are categories of models $\Mod(T)$ of suitable theories $T$ 
in a language $L_{\kappa\lambda}(\Sigma)$ where $\Sigma$ is a signature. Morphisms in $\Mod(T)$ are homomorphisms.  
Makkai and Par\'e found purely categorical characterizations of categories of this form and showed, among many other facts, 
that for every theory $T$ of $L_{\kappa\lambda}(\Sigma)$, $\Elem(T)$, the category of $T$-models together with 
$L_{\kappa\lambda}$-elementary embeddings forms an accessible category.  Abstract elementary classes were introduced by 
S.~Shelah with a similar aim -- to formalize properties of models of infinitary first-order theories of finitary signatures 
(see \cite{B}), especially the cardinalities of non-isomorphic models, without explicit reference to syntax. There is 
a visible connection of abstract elementary classes with accessible categories; we realized this in 2007.  Around the same 
time, M.~Lieberman arrived at very similar results, characterizing abstract elementary classes as suitable subcategories 
of categories $\Emb(\Sigma)$ of $\Sigma$-structures and submodel embeddings where $\Sigma$ is finitary; see \cite{L} 
and \cite{L1}. The role of this ambient category can also be played by $\Str(\Sigma)$, the category of $\Sigma$-structures 
and homomorphisms, and by other finitely accessible categories.

Unlike abstract elementary classes, accessible categories are not equi\-pped with canonical `underlying sets'. Nonetheless, 
a good substitute for `size of the model', namely, the \emph{presentability rank} of an object, can be expressed purely 
in the language of category theory, i.e.\ in terms of objects (models)\ and morphisms (embeddings).  In this sense, accessible 
categories take an extreme `signature-free' and `elements-free' view of abstract elementary classes (AEC).  From this point 
of view, Shelah's Categoricity Conjecture, the driving force of AEC (see \cite{B}), turns on the subtle interaction between 
ranks of objects and directed colimits in accessible categories.

Abstract elementary classes turn out to be placed between $\Elem(T)$ where $T$ is a theory of $L_{\kappa\omega}$ and
general accessible categories with directed colimits whose morphisms are monomorphisms.  

\section {Accessible categories}
In order to define accessible categories one just needs the concept of a $\lambda$-directed colimit where
$\lambda$ is a regular cardinal number. This is a colimit over a diagram $D:\cd\to\ck$ where $\cd$ is
$\lambda$-directed poset (taken as a category). An object $K$ of a category $\ck$ is called 
$\lambda$-\textit{presentable} if its hom-functor $\hom(K,-):\ck\to\Set$ preserves $\lambda$-directed
colimits; $\Set$ is the category of sets.

A category $\ck$ is called $\lambda$-\textit{accessible}, where $\lambda$ is a regular cardinal, provided that
\begin{enumerate}
\item[(1)] $\ck$ has $\lambda$-directed colimits,
\item[(2)] $\ck$ has a set $\ca$ of $\lambda$-presentable objects such that every object of $\ck$ is a 
$\lambda$-directed colimit of objects from $\ca$.
\end{enumerate}
A category is \textit{accessible} if it is $\lambda$-accessible for some regular cardinal $\lambda$.  

\begin{defi}\label{def2.1}
{
\em
We say that a category $\ck$ is \textit{well} $\lambda$-\textit{accessible} if it is $\mu$-accessible for each regular
cardinal $\lambda\leq\mu$. 

$\ck$ is \textit{well accessible} if it is well $\lambda$-accessible for some regular cardinal $\lambda$.
}
\end{defi}

\begin{rem}\label{rem2.2}
{
\em
(i) \textit{Locally presentable categories} are defined as cocomplete accessible categories. Following \cite{AR}, 1.20, 
each locally $\lambda$-pre\-sen\-tab\-le category is well $\lambda$-accessible.
 
(ii) Let $\lambda$ be an uncountable regular cardinal. The category $\Pos_\lambda$ of $\lambda$-directed posets 
and order embeddings is $\lambda$-accessible but it is not well accessible. In fact, $\Pos_\lambda$ is $\mu$-accessible 
if and only if $\lambda\triangleleft\mu$.  More generally, if $\lambda\triangleleft\mu$ then any $\lambda$-accessible category 
$\ck$ is $\mu$-accessible, but a $\lambda$-accessible category may fail to be $\mu$-accessible for a proper class of regular 
cardinals $\lambda < \mu$.  See \cite{MP}, 2.3.10 or \cite{AR}, 2.11.

(iii) Since $\omega\triangleleft\lambda$ for each uncountable regular cardinal $\lambda$ (see \cite{AR}, 2.13 (1)),
each finitely accessible category is well finitely accessible.  
}
\end{rem}

A \textit{signature} $\Sigma$ is a set of (infinitary) relation and function symbols. These symbols are
$S$-sorted where $S$ is a set of sorts. It is advantageous to work with many-sorted signatures but it is
easy to reduce them to single-sorted ones. One just replaces sorts by unary relation symbols and adds axioms
saying that there are disjoint and cover the underlying set of a model. Thus the underlying set of an $S$-sorted
structure $A$ is the disjoint union of underlying sets $A_s$ over all sorts $s\in S$. The category of all 
$\Sigma$-structures and homomorphisms is denoted by $\Str(\Sigma)$ and the category of all $\Sigma$-structures 
and submodel embeddings by $\Emb(\Sigma)$. Both $\Str(\Sigma)$ and $\Emb(\Sigma)$ are accessible categories, cf.\ \cite{AR}, 
5.30 and 1.70.  

A signature $\Sigma$ is \textit{finitary} if all relation and function symbols are finitary. For a finitary signature,
the category $\Str(\Sigma)$ is finitely accessible and $\Emb(\Sigma)$ is well accessible and has directed colimits.
both cases, there is a cardinal $\kappa$  such that, for each regular cardinal $\kappa\leq\mu$, a $\Sigma$-structure $K$ is 
$\mu$-presentable if and only if $|K|<\mu$. This follows from the downward L\" owenheim-Skolem theorem (\cite{MP}, 3.3.1).

\begin{exam}\label{ex2.3}
{
\em
Let $\Sigma$ be one-sorted with countably many unary relation symbols $R_n$, $n=0,1,\dots,n,\dots$. Let $A=\{a\}$ be 
the $\Sigma$-structure on a singleton set $\{a\}$ that satisfies all relations $R_n$.  Then $A$ is not finitely presentable, 
and does not have any non-trivial finitely presentable submodels.  Thus $\Emb(\Sigma)$ is not finitely accessible.  
It is, however, $\omega_1$-accessible.
}
\end{exam}

Let $\lambda$ be a cardinal and $\Sigma$ be a $\lambda$-ary signature, i.e., all relation and function symbols have arity 
smaller than $\lambda$. Given a cardinal $\kappa$, the language $L_{\kappa\lambda}(\Sigma)$ allows less than $\kappa$-ary 
conjunctions and disjunctions and less than $\lambda$-ary quantifications.  A \textit{theory} $T$ is a set of sentences 
of $L_{\kappa\lambda}(\Sigma)$. $\Mod(T)$ denotes the category of $T$-models and homomorphisms, $\Emb(T)$ the category 
of $T$-models and submodel embeddings while $\Elem(T)$ will denote the category of $T$-models and $L_{\kappa\lambda}$-elementary 
embeddings. The category $\Elem(T)$ is accessible (see \cite{AR}, 5.42).  For certain theories $T$, the category $\Mod(T)$ 
does not have $\mu$-directed colimits for any regular cardinal $\mu$ and thus fails to be accessible.

A theory $T$ is called \textit{basic} if it consists of sentences
$$
\forall(x)(\varphi(x)\Rightarrow\psi(x))
$$
where $\varphi$ and $\psi$ are positive-existential formulas and $x$ is a string of variables.  For a basic theory $T$, 
the category $\Mod(T)$ is accessible. Conversely, every accessible category is equivalent to the category of models 
and homomorphisms of a basic theory. All these facts can be found in \cite{MP} or \cite{AR}. Let us add that, prior to \cite{MP}, 
they were announced in \cite{R}.
 
A functor $F:\ck\to\cl$ is called $\lambda$-accessible if $\ck$ and $\cl$ are $\lambda$-accessible categories and $F$ preserves 
$\lambda$-directed colimits. $F$ is called accessible if it is $\lambda$-accessible for some regular cardinal $\lambda$. 
By the Uniformization theorem (see \cite{MP}, 2.5.1 or \cite{AR}, 2.19), for each accessible functor $F$ there is a regular 
cardinal $\lambda$ such that $F$ is $\lambda$-accessible and preserves $\lambda$-presentable objects; it means that if $K$ 
is $\lambda$-presentable then $F(K)$ is $\lambda$-presentable. The same is then true for each $\mu\triangleright\lambda$.  

\begin{defi}\label{def2.4}
{
\em
A functor $F:\ck\to\cl$ will be called \textit{well} $\lambda$-\textit{accessible} if $\ck$ and $\cl$ are well $\lambda$-accessible 
categories and $F$ preserves $\lambda$-directed colimits and $\mu$-presentable objects for each $\lambda\leq\mu$. $F$ is called 
\textit{well acce\-ssib\-le} if it is well $\lambda$-accessible for some regular cardinal $\lambda$. 
}
\end{defi} 

\begin{rem}\label{re2.5}
{
\em
(i) Every colimit preserving functor between cocomplete $\lambda$-accessible categories is well $\lambda$-accessible.
It immediatelly follows from the fact that, for regular cardinals $\lambda\leq\mu$, an object of a locally 
$\lambda$-presentable category is $\mu$-presentable if and only if it is a $\mu$-small colimit of $\lambda$-presentable
objects. Recall that a category is $\mu$-\textit{small} if it has less than $\mu$ morphisms.

(ii) Every finitely accessible functor is well finitely accessible. It follows from \cite{AR}, 2.13 (1) and the fact
that, for regular cardinals $\lambda\triangleleft\mu$, an object of a $\lambda$-presentable category is $\mu$-presentable 
if and only if it is a $\lambda$-directed $\mu$-small colimit of $\lambda$-presentable objects (see \cite{MP}, 2.3.11). 
}
\end{rem}

We say that a functor $F:\ck\to\cl$ \textit{reflects} $\lambda$-\textit{presentable objects} if $F(K)$ $\lambda$-presentable 
implies that $K$ is $\lambda$-presentable.

Recall that a morphism $g:B\to A$ is a split epimorphism if there exists $f:A\to B$ with $gf=\id_A$. Since, in this case, $g$
is a coequalizer of $fg$ and $\id_B$, $A$ is $\lambda$-presentable whenever $B$ is $\lambda$-presentable (see \cite{AR}, 1.16).

\begin{defi}\label{def2.6}
{
\em
We say that a functor $F:\ck\to\cl$ \textit{reflects split epimorphisms} if $f$ is a split epimorphism whenever $F(f)$ is a split
epimorphism.
 }
\end{defi} 

\begin{lemma}\label{le2.7}
Let $F:\ck\to\cl$ be a $\lambda$-accessible functor which reflects split epimorphisms. Then $F$ reflects $\lambda$-presentable 
objects. 
\end{lemma}
\begin{proof}
Let $F(K)$ be $\lambda$-presentable in $\cl$. Since $\ck$ is $\lambda$-accessible, $K$ is a $\lambda$-directed colimit
$(k_i:K_i\to K)_{i\in I}$ of $\lambda$-presentable objects. Since $F$ preserves $\lambda$-directed colimits and $F(K)$ 
is $\lambda$-presentable, there is $i\in I$ and $f:F(K)\to F(K_i)$ such that $F(k_i)f=\id_{F(K)}$. Since $F$ reflects split 
epimorphisms, there is $k_i:K_i\to K$ is a split epimorphism. Thus $K$ is $\lambda$-presentable.
\end{proof}

\begin{defi}\label{def2.8}
{
\em
A subcategory $\ck$ of a category $\cl$ will be called \textit{nearly full} if for each commutative triangle 
$$
\xymatrix@=3pc{
A \ar[rr]^{h}
\ar[dr]_{f} && C\\
& B \ar[ur]_{g}
}
$$
$h$ and $g$ in $\ck$ imply that $f$ is in $\ck$.  
}
\end{defi} 

\begin{rem}\label{re2.9}
{
\em
\cite{Ki} calls such subcategories coherent. The embedding of a nearly full subcategory $\ck$ in $\cl$ reflects split epimorphisms.
}
\end{rem}

\begin{defi}\label{def2.10}
{
\em
Let $\lambda$ be a regular cardinal.  We say that an object $K$ of a category $\ck$ has \textit{presentability rank} $\lambda$
if it is $\lambda$-presentable but not $\mu$-presentable for any regular cardinal $\mu<\lambda$.  We will write $\rank(K)=\lambda$.
}
\end{defi}

This concept was introduced by Makkai and Par\'e (see the very last line of p.29 of \cite{MP})\ under the name of 
\emph{presentability} of an object.  
 
\begin{rem}\label{re2.11}
{
\em
(1) We have mentioned that in $\Emb(\Sigma)$, $\Sigma$ finitary, there is a cardinal $\kappa$ such that, for each cardinal 
$\kappa\leq\mu$, a $\Sigma$-structure $A$ is $\mu$-presentable if and only if $|A|<\mu$.  Therefore for all large enough 
structures, $\rank(A)=|A|^+$, i.e., presentability ranks play the role of cardinalities.

(2) Assuming GCH, in any accessible category $\ck$, $\rank(K)$ is a successor cardinal for all objects $K$ (with the possible 
exception of a set of isomorphism types).  Indeed, let $\ck$ be a $\kappa$-accessible category and $\rank(K)=\lambda$ 
with $\kappa<\lambda$. $\lambda$ is a regular cardinal by definition; if $\lambda$ was a limit cardinal then (as it is 
uncountable)\ it would be a weakly inaccessible cardinal. Since GCH is assumed, $\lambda$ is inaccessible. Thus, given 
$\alpha < \kappa <\lambda$ and $\beta <\lambda$, we have $\beta^\alpha <\lambda$. Following \cite{AR} 2.13 (4), 
$\kappa\triangleleft\lambda$.  Following \cite{MP}, 2.3.11, the object $K$ can be exhibited as a $\kappa$-directed colimit 
of $\kappa$-presentable objects along a diagram of size less than $\lambda$.  Let the size of that diagram be $\nu$; then 
$K$ is $\nu^+$-presentable. That would mean $\nu^+=\lambda$, contradicting that $\lambda$ is a limit cardinal.
}
\end{rem}

\begin{propo}\label{prop2.12}
Let $F:\ck\to\cl$ be a well $\lambda$-accessible functor which reflects split epimorphisms. Then $F$ preserves presentability 
ranks $\mu$ for $\lambda<\mu$.
\end{propo}
\begin{proof}
Let $\rank(K)=\mu$, $\lambda<\mu$. Then $F(K)$ is $\mu$-presentable. Assume that $F(K)$ is $\nu$-presentable for some $\nu<\mu$. 
Without loss of generality, we can assume that $\lambda\leq\nu$. Following \ref{le2.7}, $K$ is $\nu$-presentable, which 
is a contradiction. Thus $\rank(F(K))=\mu$.
\end{proof}

\section{Accessible categories with directed colimits}
\begin{propo}\label{prop3.1}
Each accessible category with directed colimits is well accessible. 
\end{propo}
\begin{proof}
Let $\ck$ be a $\lambda$-accessible category with directed colimits and consider a regular cardinal $\lambda<\mu$. Given 
an object $K$ of $\ck$, there is a $\lambda$-directed colimit $(a_i:A_i\to K)_{i\in I}$ of $\lambda$-presentable objects $A_i$. 
Let $\hat{I}$ be the poset of all directed subsets of $I$ of cardinalities less than $\mu$ (ordered by inclusion). Since 
$\omega\triangleleft\mu$, $\hat{I}$ is $\mu$-directed (see \cite{AR}, 2.13 and 2.11). For each $M\in\hat{I}$, let $B_M$ be 
a colimit of a subdiagram indexed by $M$. Then $B_M$ is $\mu$-presentable and $K$ is a $\mu$-directed colimit of $B_M$, 
$M\in\hat{I}$ (see the proof of the last implication in \cite{AR}, 2.11).
\end{proof}

\begin{propo}\label{prop3.2}
Let $\ck$ and $\cl$ be accessible categories with directed colimits and $F:\ck\to\cl$ a functor preserving directed colimits.
Then $F$ is well accessible. 
\end{propo}
\begin{proof}
There is a regular cardinal $\lambda$ such that both $\ck$ and $\cl$ are $\lambda$-accessible and $F$ preserves 
$\lambda$-presentable objects (see \cite{AR} 2.19). Consider a regular cardinal $\lambda<\mu$ and let $K$ be a $\mu$-presentable 
object of $\ck$. Following the proof of \ref{prop3.1}, $K$ is a $\mu$-directed colimit of objects $B_M$ where each $B_M$ is 
a directed colimit of less than $\mu$ $\lambda$-presentable objects. Since $K$ is $\mu$-presentable, it is a retract of some 
$B_M$. Since $F(B_M)$ is a directed colimit of less than $\mu$ $\lambda$-presentable objects, $F(B_M)$ is $\mu$-presentable 
in $\cl$. Thus $F(K)$ is $\mu$-presentable in $\cl$ as a retract of $F(B_M)$. We have proved that $F$ is well $\lambda$-accessible.
\end{proof}

Recall that a full subcategory $\ck$ of a category $\cl$ is called accessibly embedded if there is a regular cardinal $\lambda$ 
such that $\ck$ is closed under $\lambda$-directed colimits in $\cl$.

\begin{theo}\label{th3.3}
Accessible categories with directed colimits are precisely reflective and accessibly embedded subcategories of finitely accessible
categories.
\end{theo}
\begin{proof}
Every reflective and accessibly embedded subcategory of a fi\-ni\-te\-ly accessible category $\cl$ is accessible (see \cite{AR}, 
2.53) and has directed colimits calculated as reflections of directed colimits in $\cl$. Conversely, let $\ck$ be 
a $\lambda$-accessible category with directed colimits. Then $\ck$ is equivalent to a full subcategory $\Ind_{\lambda}(\cc)$ 
of $\Set^{\cc^{\op}}$ consisting of all $\lambda$-directed colimits of hom-functors (see \cite{AR}, 2.26); $\cc$ is a small 
category. Let $\Ind(\cc)$ be the full subcategory of $\Set^{\cc^{\op}}$ consisting of all directed colimits of hom-functors. 
Then $\Ind_{\lambda}(\cc)$ is closed in $\Ind(\cc)$ under $\lambda$-directed colimits. Given an object $X$ in $\Ind(\cc)$, 
we express it as a directed colimit of hom-functors and take their colimit $F(X)$ in $\Ind_{\lambda}(\cc)$ (recall that $\ck$ 
has directed colimits). Clearly, $F(X)$ is a reflection of $X$ in $\Ind_{\lambda}(\cc)$. Thus $\Ind_{\lambda}(\cc)$ is 
a reflective subcategory of $\Ind(\cc)$.
\end{proof}

\begin{defi}\label{def3.4}
{
\em
An accessible category $\ck$ will be called $\lambda$-\textit{LS-accessible} if $\ck$ has an object of presentability rank 
$\mu^+$ for each cardinal $\mu\geq\lambda$.

$\ck$ is \textit{LS-accessible} if it is $\lambda$-LS-accessible for some cardinal $\lambda$.
}
\end{defi} 

\begin{rem}\label{re3.5}
{
\em
(1) Following \ref{re2.11} (1), $\Emb(\Sigma)$ is LS-accessible for each finitary signature $\Sigma$ and $\rank(X)=|X|^+$ 
for each $\Sigma$-structure $X$. 
 
(2) Let $T$ be a basic theory in a language $L_{\kappa\omega}(\Sigma)$. Since $\Emb(T)$ is closed under directed colimits in $\Emb(\Sigma)$,
$\rank(X)\leq|X|^+$ for each $T$-model $X$. Following the downward L\" owenheim-Skolem theorem (\cite{MP} 3.3.1) there is a cardinal $\mu$
such that each $T$-model is a $\mu^+$-directed colimit of submodels which are $T$-models and have cardinality $\mu$. Thus, for $\mu\leq|X|$, 
we have $\rank(X)=|X|^+$. Moreover, $\Emb(T)$ is LS-accessible. 
 
(3) Given an arbitrary theory $T$ a language $L_{\kappa\omega}(\Sigma)$, we can use Morley's device, to find a conservative extension $T^+$ 
of the theory $T$ such that $T^+$-homomorphisms are those $T$-homomorphisms that preserve and reflect the prescribed formulas (see \cite{MP}
3.2.8  for details). Consequently, $\Elem(T)$ is LS-accessible and $\rank(X)=|X|^+$ for each $T$-model $X$. 
}
\end{rem}

\begin{theo}\label{th3.6}
Each locally presentable category is LS-accessible. 
\end{theo}
\begin{proof}
Let $\ck$ is a locally presentable category. Then $\ck$ is locally $\lambda$-presentable for some regular cardinal $\lambda$
and, following \cite{AR} 5.30, $\ck$ is equivalent to $\Mod(T)$ for a limit theory of $L_{\lambda\lambda}(\Sigma)$ where $\Sigma$ 
is an $S$-sorted signature. Thus it suffices to prove that $\Mod(T)$ is LS-accessible.

Let $\Set^S$ denote the category of $S$-sorted sets. The category $\Set^S$ is locally finitely presentable and, given an $S$-sorted
set $X=(X_s)_{s\in S}$, $\rank(X)=|X|^+$ where $|X|$ is defined as the cardinality of the disjoint union of $X_s$, $s\in S$. Let
$U:\ck\to\Set^S$ denote the forgetful functor. The functor $U$ preserves limits and $\lambda$-directed colimits (see the proof 
of \cite{AR} 5.9). Thus $U$ has a left adjoint $F$ (\cite{AR} 1.66) which preserves $\mu$-presentable objects for each
$\mu\geq\lambda$ (\cite{AR} Ex. 1.s(1)). Thus $\rank(FX)\leq\rank(X)$ for each $X$ with $\lambda\leq\rank(X)$. Assume that
$\mu=\rank(FX)<\rank(X)$ for $X$ with $\lambda\leq\rank(X)$. Since $\Set^S$ is locally $\mu$-presentable, $X$ is a $\mu$-directed
colimit of $\mu$-presentable objects $X_i$, $i\in I$. Let 
$$
(u_i:X_i\to X)_{i\in I}
$$ 
denote a colimit cocone. Since $F$ preserves colimits, 
$$
(Fu_i:FX_i\to FX)_{i\in I}
$$ 
is a colimit cocone. Since $FX$ is $\mu$-presentable, there is $j\in I$ and $r:FX\to FX_j$ such that $F(u_j)r=\id_{FX}$. Hence $Fu_j$ 
is a split epi\-mor\-phism. We can assume, without a loss of generality, that $X_s\neq\emptyset$ implies $(X_i)_s\neq\emptyset$ 
for each $s\in S$. Then each $u_i$ is a split monomorphism and thus each $Fu_i$ is a split monomorphism. Consequently, $Fu_j$ 
is an isomorphism and $r=(Fu_j)^{-1}$. 

Let $\eta:\Id\to UF$ be the adjunction unit. Then 
$$
U(r)\eta_X:X\to UFX_j
$$ 
makes $FX_j$ free over $X$. This contradicts to the fact that each endomorphism $f:FX_j\to FX_j$ with $U(f)\eta_{X_j}=\eta_{X_j}$ 
is equal to $\id_{FX_j}$. Indeed, since
$$
U(r)\eta_Xu_j=U(r)UF(u_j)\eta_{X_j}=\eta_{X_j},
$$ 
the image of $X$ in $U(r)\eta_X$ extends $X_j$ to free generators of $UFX_j$. Then, by interchanging two free generators not belonging to $X_j$,
we get an $f$ above with $f\neq\id_{FX_j}$.
\end{proof}

\section {Abstract elementary classes}
A class $\ck$ of $\Sigma$-structures will be always considered as the full subcategory of $\Str(\Sigma)$.
The following definition introduces abstract elementary classes using the language of category theory.
$|A|$ denotes the cardinality of the underlying set of the $\Sigma$-structure $A$.

\begin{defi}\label{def4.1}
{
\em
Let $\Sigma$ be a finitary signature, $\ck$ a class of $\Sigma$-structures and $\cm$ a class of homomorphisms. 
Then $(\ck,\cm)$ is an \textit{abstract elementary class} if the following conditions are satisfied:
\begin{enumerate}
\item[(A1)] each homomorphism $h:A\to B$ in $\cm$ is an embedding of $A\in\ck$ to $B\in\ck$ as a submodel,
\item[(A2)] $\ck$-objects together with $\cm$-homomorphisms form a category $\ck_\cm$,
\item[(A3)] $\ck_\cm$ is closed in $\Str(\Sigma)$ under directed colimits,
\item[(A4)] given homomorphisms $g:A\to B$ and $h:B\to C$ with $hg,h\in\cm$ then $g\in\cm$, and
\item[(A5)] there is a cardinal $\lambda$ such that if $f:A\to B$ is a submodel embedding with $B\in\ck$
then there is $h:A'\to B$ in $\cm$ such that $f$ factorizes through $h$ and $|A'|\leq |A|+\lambda$.
\end{enumerate}
}
\end{defi}

The standard formulation (see \cite{B}) uses colimits of continuous chains in (A3). But it is well known that it implies
that it is true for all directed colimits (see \cite{B}, or \cite{AR} 1.7; continuous chains are called smooth there). 
The standard formulation also includes that both $\ck$ and $\cm$ are closed under isomorphisms. This means that, given $K$ 
in $\ck$ and an isomorphism $h:K\to X$ with $S$ in $\Str(\Sigma)$ then $X$ belongs to $\ck$ and $h$ belongs to $\cm$; 
the first property means that $\ck$ is \textit{replete} in $\Str(\Sigma)$. But this closedness under isomorphisms 
is a consequence of (A3). In (A4), we do not need to assume that $g$ is a submodel embedding
because it follows from $hg$ being a submodel embedding. Let us recall that submodel embeddings are precisely
strong homomorphisms in $\Str(\Sigma)$. This is the standard category theoretic terminology and our reason
for avoiding to call $\cm$-morphisms strong embeddings as one does in the abstract elementary classes literature.

We allow many-sorted signatures in \ref{def4.1}. Since each many-sorted signature can be made single-sorted, this does not change the concept 
of an abstract elementary class. The following result was also observed in \cite{L} Theorem 5.9 (see \cite{L1} Theorem 4.1, as well).
There is a related work \cite{Ki}.
 
\begin{propo}\label{prop4.2}
Let $(\ck,\cm)$ be an abstract elementary class. Then $\ck_\cm$ is an LS-accessible category with directed colimits.
\end{propo}
\begin{proof}
We observed in \ref{re3.5} that $\Emb(\Sigma)$ is an LS-accessible category with directed colimits. There is a regular cardinal
$\kappa$ greater than $\lambda$ from (A5) such that $\Emb(\Sigma)$ is $\kappa$-accessible and $\kappa$-presentable objects 
in $\Emb(\Sigma)$ are precisely $\Sigma$-structures $A$ such that $|A|<\kappa$. (A5) then implies that each object of $\ck_\cm$ 
is a $\kappa$-directed colimit of $\kappa$-presentable objects in $\ck_\cm$. Thus $\ck_\cm$ is $\kappa$-accessible. Moreover,
$\ck_\cm$ has an exactly $\mu$-presentable object  for each regular cardinal $\kappa\leq\mu$. Thus $\ck_\cm$ is LS-accessible.
Following (A5), the embedding of $\ck_\cm$ into $\Emb(\Sigma)$ preserves directed colimits and preserves exactly $\mu$-presentable
objects for each regular cardinal $\kappa\leq\mu$.
\end{proof}

\begin{rem}\label{rem4.3}
{
\em
(1) Following the proof above, we have $\rank(A)=|A|^+$ for each $A$ in $\ck$ with $|A|\geq\lambda$.

(2) $\ck_\cm$ is well accessible (see \ref{prop3.1}).
}
\end{rem}

The following result is closely related to Claim 4.9 in \cite{L1}. We just avoid talking about structures and signatures.

\begin{theo}\label{th4.4} A category $\cl$ is equivalent to $\ck_\cm$ for an abstract elementary class $(\ck,\cm)$ if and only if
it is an accessible category with directed colimits, whose morphisms are monomorphisms and which admits a nearly full embedding
into a finitely accessible category preserving directed colimits and monomorphisms.
\end{theo}
\begin{proof}
Necessity is evident (using \ref{prop4.2}). Consider a category $\cl$ satisfying the conditions above for a finitely accessible category
$\ch$. Let $\ca$ be a representative set of finitely presentable objects in $\ch$. Consider the canonical functor
$$
E:\ch\to\Set^{\ca^{\op}}.
$$
(see \cite{AR}, 1.25). Since objects of $Set^{\ca^{\op}}$ can be viewed as many-sorted unary algebras, $Set^{\ca^{\op}}$ is 
a full subcategory of $\Str(\Sigma)$ for a finitary signature $\Sigma$. Since $E$ is a full embedding, we get a nearly full embedding
$$
\cl\to\Str(\Sigma).
$$
Since subalgebras are submodels, condition (A1) is satisfied. The same is evidently true for (A2)-(A4). Following \ref{prop3.2} 
and \ref{prop3.1}, $G$ is well $\lambda$-accessible for some regular cardinal $\lambda$. Moreover, we can assume that an object
$X$ in $\Emb(\Sigma)$ is $\mu$-presentable if and only if $|X|<\mu$ for each $\mu\geq\lambda$. Let $f:A\to G(B)$ be a submodel embedding. 
Then $G(B)$ is $\mu$-presentable in $\Emb(\Sigma)$ for some regular cardinal $\mu\geq\lambda$. Since $B$ is a $\mu$-directed colimit 
of $\mu$-presentable objects, there is $h:A'\to B$ in $\ck$ such that $A'$ is $\mu$-presentable in $\ck$ and $f$ factorizes 
through $h$. Since $G(A')$ is $\mu$-presentable in $\Emb(\Sigma)$, $|A'|<\mu < |A|+\lambda$. Thus (A5) is satisfied.
\end{proof}

\begin{exam}\label{ex4.5}
{
\em
(1) Let $T$ be a theory in a language $L_{\kappa\omega}(\Sigma)$. Then $\ck=\Mod(T)$ is an abstract elementary
class where $\cm$ consists of $L_{\kappa\omega}$-elementary embeddings. Thus $\ck_\cm=\Elem(T)$. 

(2) Let $T$ be a basic theory in a language $L_{\kappa\omega}(\Sigma)$. Then $\ck=\Mod(T)$ is an abstract elementary
class where $\cm$ consists of submodel embeddings. Thus $\ck_\cm=\Emb(T)$.  

(3) Let $\Sigma$ be a single-sorted signature containing just a unary relation symbol $P$. A simple example of an abstract 
elementary class $\ck$ in $\Sigma$ which is not closed under $L_{\infty\omega}$-elementary equivalence (thus not axiomatizable
in any $L_{\kappa\omega}(\Sigma)$) is given in \cite{K}, 2.10. One takes $\Sigma$-structures $K$ such that $P_K$
is countable and the complement of $P_K$ is infinite; $\cm$ consists of submodel embeddings which are identities on $P$. 
But, $\ck_\cm$ is isomorphic to the category of infinite sets and monomorphisms. The later category is $L_{\omega\omega}$ 
axiomatizable (in the empty signature). Thus, although that $(\ck,\cm)$ is not axiomatizable in the signature $\Sigma$, $\ck_\cm$
is equivalent (as a category) to an axiomatizable one.
}
\end{exam}

\begin{rem}\label{rem4.6}
{
\em
(1) Shelah's Presentation Theorem (see \cite{B}, 4.15) implies that, for each abstract elementary class $(\ck,\cm)$
in a signature $\Sigma$, there is a signature $\Sigma\subseteq\Sigma'$ and a basic theory $T$ of 
$L_{\kappa\omega}(\Sigma')$ such that $\ck$ is the full image of the reduct
$$
R:\Emb(T)\to\Str(\Sigma).
$$
The reason is that omitting a type of an element means to negate the conjunction of all atomic formulas from
this type, which is a basic formula. Moreover, $\cm$ is the full image in $R$ of $\Sigma'$-submodel embeddings.  
Since $R$ is finitely accessible, each abstract elementary class is a finitely accessible full image of that 
from \ref{ex4.5} (2).

(2) One cannot expect that abstract elementary classes are closed under finitely accessible full images. 
Let $\Sigma$ contain a (single-sorted) unary relation symbol $U$ and $\ck$ consist of all $\Sigma$-structures $A$ such that 
$|A|=|U_A|$. Since $\ck$ is not closed under directed colimits of submodel embeddings, $(\ck,\cm)$ is not an abstract elementary 
class for $\cm$ consisting of submodel embeddings. But $\ck_\cm$ form the full image of 
$$
R:\Emb(T)\to\Str(\Sigma)
$$ 
where $\Sigma'$ also contains a binary operation symbol $f$ and $T$ consists of formulas
$$
(\forall x)U(f(x))
$$
and
$$
(\forall x,y)(f(x)=f(y)\to x=y).
$$
Since $R$ is finitely accessible, $\Emb(T)$ is an abstract elementary class with a finitely accessible full image which 
is not an abstract elementary class.
}
\end{rem}

\section{Categoricity}
\begin{defi}\label{def5.1}
{
\em
Let $\lambda$ be an infinite cardinal. A category $\ck$ is called $\lambda$-\textit{categorical} it it has,
up to isomorphism, precisely one object of the presentability rank $\lambda^+$.
}
\end{defi} 

\begin{rem}\label{rem5.2}
{
\em
Following \ref{rem4.3} (1), the definition \ref{def5.1} (suggested in \cite{R2}) is in accordance with its model theoretic meaning 
for abstract elementary classes.
}
\end{rem}

Shelah's Categoricity Conjecture claims that for every abstract elementary class $(\ck,\cm)$ there is a cardinal
$\kappa$ such that if $\ck_\cm$ is $\kappa$-categorical then it is $\lambda$-categorical for each $\kappa<\lambda$.  
Following the remark above, this is a statement about the category $\ck_\cm$ -- it does not depend on its signature presentation. 

\begin{exam}\label{ex5.3}
{
\em
Consider a one-sorted signature $\Sigma$ given by a sort $s$ and an $\omega$-ary function symbol $f:s^\omega\to s$. Let $T$ be 
the $L_{\omega_1\omega_1}(\Sigma)$ theory saying that $f$ is a bijection.  Thus $T$ consists of the formula
$$
(\forall y)(\exists !x)(f(x)=y)
$$
where $x$ is an $\omega$-string of variables, and $\exists !$ denotes unique existence. By general facts 
about categories of models of $\exists !$-sentences (called \textit{li\-mit theories}), the category $\Mod(T)$ is locally 
$\omega_1$-presentable; see \cite{AR}, 5.30. Hence, in particular, it is accessible and has directed colimits.
 
If $\mu$ has cofinality $\omega$ then 
$$
\mu^\omega > \mu
$$
cf.\ \cite{J}, Corollary 4 of Theorem 17.  Thus $T$ does not have models of cardinalities of cofinality $\omega$. Since
$\mu^\omega=\mu$ whenever the cardinal $\mu$ is of the form $\nu^\omega$, $\Mod(T)$ is a large category. Thus there is a proper
class of cardinalities in which $T$ has a model and a proper class of cardinalities in which $T$ does not have a model. But,
following \ref{th3.6}, $\Mod(T)$ is LS-accessible. The point is that presentation ranks and cardinalities will never start
to coincide. 
}
\end{exam}

\begin{exam}\label{ex5.4}
{
\em
Let $\ck$ be an accessible category which is not LS-acce\-ssib\-le.  Let $\cl=\ck\sqcup\Set$, the disjoint union of $\ck$ and the category of sets.  
$\cl$ is an accessible category and there is a proper class of cardinals $\lambda$ such that $\cl$ is $\lambda$-categorical and, at the same time, 
a proper class of cardinals $\lambda$ such that $\cl$ is not $\lambda$-categorical.
}
\end{exam}

But we do not know any example of an accessible category with directed colimits which is not LS-accessible. In the same time, we do not
know any example of an accessible category with directed colimits, whose morphisms are monomorphisms, and which is not equivalent to
$\ck_\cm$ for any abstract elementary class $(\ck,\cm)$. We expect, at least in the second case, that such an example exists. Difficulty
in producing it is well illustrated by the next example.  
 
\begin{exam}\label{ex5.5}
{
\em
Let $\ck$ be the category of commutative unital $C^\ast$-al\-geb\-ras. Gel\-fand duality says that $\ck$ is isomorphic
to the dual of the category of compact Hausdorff topological spaces. $\ck$ is a locally $\omega_1$-pre\-sen\-tab\-le category.
Thus it is an accessible category with directed co\-li\-mits. There is not known whether $\ck$ is equivalent to $\Mod(T)$ where
$T$ is a theory of $L_{\kappa\omega}$. P. Bankstone \cite{Bn} asked whether $\ck$ is equi\-va\-lent to $\Mod(T)$ where $T$ is 
an $L_{\omega\omega}$ theory of a signature $\Sigma$ such that $\Mod(T)$ is closed in $\Str(\Sigma)$ under products.
A negative answer was independently given in \cite{Ba} and \cite{R3}.  
}
\end{exam}

\end{document}